# Visual Cue Enhancement and Dual Low-Rank Adaptation for Efficient Visual Instruction Fine-Tuning

### v2

@@ Abstract
Parameter-efficient fine-tuning multimodal large language models (MLLMs) presents significant challenges, including a reliance on high-level visual features that limits fine-grained detail comprehension, and data conflicts that arise from task complexity. To address these issues, we propose an efficient fine-tuning framework with two novel approaches: Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA). VCE enhances the vision projector by integrating multi-level visual cues, improving the model’s ability to capture fine-grained visual features. Dual-LoRA introduces a dual low-rank structure for instruction tuning, decoupling learning into skill and task spaces to enable precise control and efficient adaptation across diverse tasks. Our method simplifies implementation, enhances visual comprehension, and improves adaptability. Experiments on both downstream tasks and general benchmarks demonstrate the effectiveness of our proposed approach.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across a wide range of tasks, showcasing versatility and effectiveness in diverse applications. Existing MLLMs [14, 1, 15] typically integrate pretrained Large Language Models (LLMs) [26, 5] with pretrained vision encoders [22, 16], using a vision projector to connect visual features to the LLMs.
The standard training process for MLLMs generally involves two stages: (1) a vision projector pretraining stage, where only the vision projector is trained to align visual features for the LLMs, and (2) an instruction tuning stage, where both the vision projector and the LLM are trained to perform visual instruction tasks. To reduce the extensive number of parameters in LLM tuning, low-rank adaptation (LoRA) [8] was proposed by injecting lightweight adapters into pretrained LLMs for efficient fine-tuning. The vision projector pretraining aligns visual and language feature spaces, while instruction tuning equips MLLMs to generate responses aligned with human instructions. The goal of these two stages is to equip MLLMs with sufficient visual cues and to ensure accurate instruction-based responses.
Despite the great success, existing MLLMs have two main limitations. On the one hand, in the vision projector pretraining stage, current methods primarily rely on high-level visual features, often overlooking low-level and fine-grained details, which limits visual comprehension. For example, most works[15, 2] use only the high-level semantic feature map for vision token projection, while some [24, 31] leverage multi-layer visual features yet still miss finer details. On the other hand, in the instruction-tuning stage, as downstream tasks grow in diversity and complexity, data conflicts in LoRA-based instruction tuning become more prominent  [4, 9]. To mitigate this, recent studies [4, 18] incorporate the Mixture of Experts (MoE) paradigm into LoRA modules, embedding multiple LoRA units within a single linear layer so that each unit can leverage its specific strengths. To better accommodate varied tasks, LoRA experts with different levels of granularity [10] or customized routing strategies [18] have been introduced. However, this complex design not only increases implementation complexity but may also extend the time required for training and inference. We thus raise the question: Is there a simple and unified approach to flexibly represent any group of LoRA experts?
Figure 1: The two training stages of our framework. In the Vision Projector Pretraining stage, we incorporate the proposed Visual Cue Enhancement module (VCE) to enhance the final vision feature map by capturing detailed local information from multi-level visual feature maps. In the Visual Instruction Fine-Tuning stage, we use the proposed Dual Low-Rank Adaptation (Dual-LoRA) as the LLM adapter to learn task-specific knowledge while mitigating data conflicts, which is set trainable alone with VCE and vision projector.
In this paper, we address existing limitations by proposing Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA) to improve the vision projector pretraining stage and the instruction finetuning stage respectively. An overview of our overall training pipeline is shown in Figure 1. For the VCE module, we extend feature capture to multi-level vision feature maps instead of relying solely on high-level semantic feature maps. Specifically, for each patch in the high-level feature map, we extract neighboring patches from the middle-layer feature maps and combine them with the original patch to create a locally enhanced feature patch. This approach enhances local visual cues while maintaining computational efficiency. To address data conflicts during MLLM instruction tuning, we introduce Dual-LoRA, a unified adaptation module that decouples learning into a skill low-rank space and a task-activation low-rank space. The skill space captures specific knowledge in downstream tasks, while the task space activates prior knowledge based on specific instructions. These two spaces are then jointly mapped into a cohesive single space to generate features for the subsequent stages. This separation of skill and task spaces within a single LoRA module enables fine-grained control over skill knowledge activation, allowing for dense LoRA rank learning of skill knowledge while supporting sparse, task-specific responses.
Experiments conducted on downstream tasks within the food domain, as well as on several general MLLM benchmarks, highlight the effectiveness of our proposed method. Our contributions can be summarized as follows:
@@
  • We introduce a lightweight Visual Cue Enhancement (VCE) module to enrich fine-grained visual features for vision projector pretraining.
@@
  • We propose Dual Low-Rank Adaptation (Dual-LoRA), a simple yet efficient approach that mitigates data conflicts for effective Visual Instruction Fine-Tuning.
  • Experimental results on both downstream tasks and general MLLM benchmarks validate the effectiveness of our proposed method.
Figure 2: Illustration of our proposed components. The Visual Cue Enhancement (VCE) module uses the projected vision feature map $f^{\textbf{a}}$ selected during pretraining as the anchor feature, leveraging nearby patches in multi-layer feature maps $[f^{\textbf{ref}}]$ to enrich the local visual cues of $f^{\textbf{a}}$. The Dual Low-Rank Adaptation (Dual-LoRA) decomposes the input feature into skill space and task space via $S$ and $T$, which are then multiplied element-wise, projected back to the same dimension as $x$ via $B$, added to the features obtained from the pretrained weights $W$, and passed to the next layer.
@@ 2 Related Works
Multi-Modal Large Language Models (MLLMs) integrate pretrained vision encoders with pretrained large language models (LLMs), showcasing remarkable instruction-following and generalization capabilities. Early models like Flamingo and BLIP-2 successfully adapted LLMs for visual tasks, demonstrating strong zero-shot generalization and in-context learning abilities. Visual instruction tuning, a key technique in recent studies [14], has further enhanced MLLMs by using vision-language (VL) datasets and improving visual instruction-following data. Additionally, some research approaches vision as a language, employing textual annotations (e.g., location descriptions of bounding boxes) to train MLLMs effectively.
@@
Enhancing Vision Feature Representation in MLLMs primarily involves vision encoder ensembling, resolution enhancement, and multi-level feature map fusion. For vision encoder ensembling, MMVP [25] utilizes a Mixture of Features (MoF) strategy to integrate image features from CLIP-ViT [22] and DINOv2 [20]. Similarly, MouSi [7] uses an ensemble technique to leverage the strengths of individual vision encoders, introducing a fusion network to unify outputs from different encoders, such as CLIP, DINOv2, and SAM. DeepSeekVL [19] adopts a hybrid vision encoder design, encoding images by combining SigLIP-L for low-resolution inputs and SAM-B for high-resolution inputs. LLaVA-UHD [29] organizes input images for efficient and scalable encoding, utilizing a compression module to further condense image tokens from visual encoders. Some approaches focus on fusing multi-level feature maps. For example, Cambrian-1 [24] introduces the Spatial Vision Aggregator, a dynamic, spatially-aware connector that integrates high-resolution visual features with LLMs while minimizing token count. While most of these methods require full fine-tuning of the LLM, we propose a lightweight visual cue enhancer for more efficient fine-tuning.
Low-Rank Adaptation (LoRA) [8] introduces a method that freezes pretrained model weights and injects trainable low-rank decomposition matrices into the linear layers of the Transformer architecture, significantly reducing the number of trainable parameters required for downstream tasks. Mixture of Experts in LoRA (MoE-LoRA) has gained popularity for addressing data conflicts in downstream task fine-tuning. MoE-LoRA typically employs three types of activation: sparse activation [4], where only the top-$k$ experts are activated; dense activation [27], where all experts are activated and re-weighted by a router; and rectified activation [10], where experts are dynamically activated and optimized to be sparse. For instance, LLaVA-MoLE selects only the top-1 expert for specific tasks. In dense activation, MoLE [27] treats each trained LoRA module at each layer as an individual expert, integrating learnable gating functions at each layer to control hierarchical weights. In rectified activation, RoDE [10] designs diverse experts and employs a rectified linear unit to optimize the router for sparse expert activation. These methods involve complex expert group configurations or routing strategies. In contrast, our approach offers a simple yet efficient adapter that effectively addresses data conflicts without requiring a complicated design.
Figure 3: Illustration of (a) VCE, (b) typical Mixture of Experts strategies for addressing data conflicts, and (c) Dual-LoRA. Mixture of Experts strategies include: Sparse activation [4], where only a subset of experts is activated; dense activation [27], where all experts contribute equally; and rectified activation [10], where all experts contribute with selective rectification.
@@ 3.1 Preliminaries
where $z_{i}$ is the output feature passed to the next layer, and $\alpha$ is a scaling factor. For convenience, we omit the indices $i$ in the following. With LoRA, the trainable parameters of $F$ are reduced to $\{\theta_{P},A,B\}$.
@@ 3.2 Framework Overview
We propose a two-stage approach for efficient Visual Instruction Tuning: (1) Enhanced Vision Projector Pretraining and (2) Dual Low-Rank Adaptation (Dual-LoRA) Visual Instruction Fine-Tuning, as illustrated in Figure 1. Specifically, to enable comprehensive vision information processing and effective execution of instruction-following downstream tasks, we propose two main components: Visual Cue Enhancement (VCE) to capture enriched visual cues by incorporating multi-level visual features, and Dual-LoRA to mitigate data conflicts by decoupling visual instruction tuning into skill-task dual low-rank space learning. The complete framework pipeline is illustrated in Figure 2, with further details provided in the following sections.
@@ 3.3 Multi-level Local Visual Cue Enhancement
Typical vision projectors primarily rely on high-level vision feature maps to align vision and language modalities. For example, LLaVA [14] and Qwen-VL [1] use the penultimate layer output of ViT, which captures high-level, language-aligned semantic features but may overlook finer, detail-oriented visual cues.
To address these limitations, we propose a lightweight Vision Cues Enhancement (VCE) module. This module uses the final vision feature map as an anchor feature and applies a single deformable cross-attention mechanism to neighboring patches across other layer feature maps for each patch of the anchor feature. An intuitive illustration is shown in Figure 3 (a).
Denote $F=[f]$ as the multi-level feature maps produced by the vision encoder $V$. We refer to the high-level feature map used for projection into vision tokens as the anchor feature $f^{\text{a}}$, while several intermediate layers are designated as reference feature maps $[f^{\text{ref}}]$. The vision encoder processes the input image $I$ through multiple blocks to extract these multi-level features $F$, formulated as:
[…]
where select-index specifies the index of reference feature maps.
To enhance the vision cues, we integrate local visual details from $[f^{\text{ref}}]$ into $f^{\text{a}}$. We utilize deformable attention [28] to efficiently compute cross-attention across multi-level feature maps, applied between $[f^{\text{ref}}]$ and $f^{\text{a}}$:
$f^{*}=\text{deform-attn}(f^{\text{a}},[f^{\text{ref}}];\theta_{\text{deform-attn}}),$ (6)
[…]
After obtaining the anchor feature $f^{\text{a}}$ and the enhanced visual cue $f^{*}$, we add them up and apply normalization to produce the final enhanced vision feature map $f_{\text{vision}}$. This feature map is subsequently projected into vision tokens $t_{\text{vision}}$ through the vision feature projector $P$:
$f_{\text{vision}}$ $=\text{norm}(f^{\text{a}}+f^{*}),$ (7)
$t_{\text{vision}}$ $=P(f_{\text{visual}};\theta_{P}).$ (8)
The enhanced vision token $t_{\text{vision}}$ is then concatenated with text tokens and fed into the large language modal model $G$ for further processing. The structure of the adaptive $G$ is detailed in the next section.
@@ 3.4 Dual Low Rank Adaptation
LoRA assumes that in large pre-trained models, the model parameters have a high degree of redundancy, with a low-rank structure that can be effectively represented through low-rank matrix approximation. However, as downstream tasks become more complex or diverse, data conflicts can arise due to the limited representational capacity of the low-rank structure [9, 4]. Some methods integrate Mixture of Experts into LoRA to better adapt to diverse downstream tasks. They propose expert modules with different granularities [10] and use various routing strategies [4, 27, 18, 10].
In contrast, we propose Dual Low-Rank Adaptation (Dual-LoRA), a unified approach that decouples adaptation into a task-skill dual low-rank structure, removing the need for complex expert splitting or intricate routing strategies. Figure 3 (b) and (c) illustrate the differences between our method and other strategies for mitigating data conflicts.

### v3

# From Holistic to Localized: Local Enhanced Adapters for Efficient Visual Instruction Fine-Tuning

@@ Abstract
Efficient Visual Instruction Fine-Tuning (EVIT) seeks to adapt Multimodal Large Language Models (MLLMs) to downstream tasks with minimal computational overhead. However, as task diversity and complexity increase, EVIT faces significant challenges in resolving data conflicts. To address this limitation, we propose the Dual Low-Rank Adaptation (Dual-LoRA), a holistic-to-local framework that enhances the adapter’s capacity to address data conflict through dual structural optimization. Specifically, we utilize two subspaces: a skill space for stable, holistic knowledge retention, and a rank-rectified task space that locally activates the holistic knowledge. Additionally, we introduce Visual Cue Enhancement (VCE), a multi-level local feature aggregation module designed to enrich the vision-language projection with local details. Our approach is both memory- and time-efficient, requiring only 1.16$\times$ the inference time of the standard LoRA method (with injection into the query and value projection layers), and just 73% of the inference time of a 4-expert LoRA-MoE. Extensive experiments on various downstream tasks and general MLLM benchmarks validate the effectiveness of our proposed methods. Our project page are publicly available at https://github.com/pengkun-jiao/Dual-LoRA
Figure 1: The motivation behind the proposed approach. (a) Humans perceive visual signals from both holistic and localized perspectives, considering the overall image structure and local details. (b) Humans acquire holistic knowledge, with specific parts (localized) applied to particular tasks. For example, one might first acquire holistic knowledge of cooking and then apply specific knowledge, e.g. ingredient recognition, to source ingredients.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated impressive performances across a wide range of tasks, showcasing versatility and effectiveness in diverse applications [16, 15, 14]. Existing MLLMs [20, 3, 21] typically integrate a pretrained Large Language Models (LLM) [32, 6] with a pretrained vision encoder [28, 22], using a vision projector to connect the vision encoder and LLM.
Vision Instruction Tuning [20] enables the practical application of MLLMs in downstream tasks. However, full fine-tuning MLLMs can be highly resource-intensive. Efficient Vision Instruction Tuning (EVIT) addresses this challenge by tuning only the vision feature projector while integrating a small number of additional trainable parameters (adapters) into the LLM [8]. For example, Low-Rank Adaptation (LoRA) [11] leverages low-rank approximations to replace full fine-tuning of LMM, significantly reducing computational overhead.
Figure 2: Comparison between the proposed Dual-LoRA and existing methods. (a) Mainstream MLLMs, e.g. LLaVA [20] and Qwen-VL [3], project the high-level visual feature map. (b) Our Visual Cue Enhancement module enhances high-level features by aggregating local information from multi-level feature maps. LoRA-MoE methods mitigate data conflicts by enabling localized responses, i.e., experts activation, tailored to several activation strategies: (c) Sparse Activation[5], where only the top-k experts are activated; (d) Dense Activation[33], where all experts are activated; and Rectified Activation [13], where multiple heterogeneous experts are dynamically activated. In contrast, our (f) Dual Low-Rank Adaptation rectifies a holistic knowledge (skill) space with an additional task space, which is fully differentiable, capable of learning any local response, and more structurally efficient.
However, existing EVIT methods face a critical limitation: as downstream tasks grow more diverse and complex, data conflicts in LoRA-based instruction tuning become increasingly pronounced [5, 12], ultimately impairing performance. An intriguing observation is that when tuning for a variety of food-related tasks, LoRA fine-tuning can result in knowledge-inconsistent answers. For example, some ingredients may be misaligned between food ingredient recognition and recipe generation, as illustrated in Fig. 4. To address this, recent studies suggest integrating LoRA within a Mixture of Experts (LoRA-MoE) paradigm [5, 23, 33]. LoRA-MoE adapts to diverse tasks by activating specific experts, which can be viewed as a process of localized adaptation, effectively mitigating data conflicts. However, the LoRA-MoE method requires a complex design to balance activation strategies, trainable parameters, and task complexity [13, 12]. Moreover, it incurs significantly higher time consumption compared to vanilla LoRA when multiple LoRA experts are activated.
To address the high time consumption of experts polling in LoRA-MoE, we propose a structurally efficient dual-space adaptation design that transitions from holistic domain knowledge learning to localized, task-specific adaptation. Additionally, we emphasize the role of localized adaptation in visual perception. A high-level overview of our motivation is illustrated in Figure 1. This design embodies a cognitively inspired learning paradigm: it first consolidates generalized skill representations and then dynamically routes task-critical knowledge components during inference, mirroring the transition in human cognition from holistic to localized processing. Specifically, we propose Dual Low-Rank Adaptation (Dual-LoRA), which strategically decouples adaptation learning into: 1) skill space learning for holistic domain knowledge acquisition, and 2) task space learning for instruction-conditioned task adaptation. The skill space is parameterized using a low-rank matrix optimized stably for cross-task knowledge consolidation, while the task space employs a rank-rectified matrix that dynamically modulates the skill space for localized task-specific adaptation. To enhance local details for vision feature projection, we further propose Visual Cue Enhancement (VCE), which enhances high-level visual features with local information aggregated from multiple feature maps. VCE remains computationally efficient by focusing on a limited number of local neighboring regions within each feature map. Figure 2 illustrates the structurally efficient design of our approach.
Our VCE module is only 5.53 MB in size, and Dual-LoRA outperforms LoRA-MoE methods while using the same number of trainable parameters. The full approach requires just 1.16$\times$ the inference time of vanilla LoRA, and only 73% of the inference time of 4-experts LoRA-MoE methods. Extensive experiments across a range of downstream tasks—including UniFood, ScienceQA, Flickr30k, and general MLLM benchmarks—demonstrate the effectiveness of our proposed methods. Our main contributions can be summarized as follows:
  • We design a time- and memory-efficient holistic-to-localized approach for EVIT. Extensive experiments demonstrate the effectiveness of our proposed methods.
@@
  • We introduce a lightweight Visual Cue Enhancement (VCE) module to enrich fine-grained visual features for efficient visual instruction fine-tuning.
@@
  • We propose Dual Low-Rank Adaptation (Dual-LoRA), an efficient approach that mitigates data conflicts for instruction fine-tuning.
@@ 2 Related Works
Multimodal Large Language Models (MLLMs) harness the powerful reasoning capabilities of Large Language Models (LLMs) [32], driving significant advancements in vision tasks. Early works, such as Flamingo [1], resample visual features and integrate them into attention-based adapter layers within the LLM. BLIP-2 [18] introduces a Q-Former, which simultaneously handles cross-modal representation learning and generative learning. More recently, Vision-Language Instruction Tuning approaches, e.g. Instruct-BLIP [7], LLaVA [20], and Mini-GPT4 [37], have curated large-scale, high-quality multimodal instruction data to enhance MLLMs’ ability to follow instructions. However, fully tuning MLLMs is highly resource-intensive due to the vast number of parameters in LLMs, which hinders the application of MLLMs to downstream tasks.
Low-Rank Adaptation (LoRA)[11] adapts the model by introducing low-rank matrices while keeping the original model parameters unchanged. Specifically, LoRA adds low-rank matrices to certain layers of the model (e.g., attention layers, feed-forward layers, etc.), enabling the model to fine-tune more efficiently for specific tasks. Mixture of Experts in LoRA (LoRA-MoE)[5, 13, 33] has gained popularity for addressing data conflicts during downstream task fine-tuning. LoRA-MoE methods can be grouped by the expert activation techniques: sparse activation [5], where only the top-k experts are activated; dense activation [33], where all experts are activated and re-weighted by a router; and rectified activation [13], where experts are dynamically activated and optimized for sparsity. RoDE [13] further introduces heterogeneous experts, where different sets of parameters are used for tasks of varying complexity. Figure 2 compares LoRA-MoE methods with our approach. LoRA-MoE activates multiple LoRA adapters, increasing inference time, whereas our method employs a single Dual-LoRA for greater inference efficiency.
@@
Enhancing Vision Feature Representation in MLLMs primarily involves vision encoder ensembling, resolution enhancement, and multi-level feature map fusion. For vision encoder ensembling, MMVP [31] utilizes a Mixture of Features (MoF) strategy to integrate image features from CLIP-ViT [28] and DINOv2 [25]. Similarly, MouSi [9] uses an ensemble technique to leverage the strengths of individual vision encoders, introducing a fusion network to unify outputs from different encoders, such as CLIP, DINOv2, and SAM. DeepSeekVL [24] adopts a hybrid vision encoder design, encoding images by combining SigLIP-L for low-resolution inputs and SAM-B for high-resolution inputs. LLaVA-UHD [35] organizes input images for efficient and scalable encoding, utilizing a compression module to further condense image tokens from visual encoders. Some approaches focus on fusing multi-level feature maps. For example, Cambrian-1 [30] introduces the Spatial Vision Aggregator, a dynamic, spatially-aware connector that integrates high-resolution visual features with LLMs while minimizing token count. While most of these methods require full fine-tuning of the LLM, our approach, in contrast, introduces a lightweight Visual Cue enhancement module for efficient fine-tuning.
Figure 3: Overview of our two key components: Visual Cue Enhancement (VCE) module enhances high-level anchor features by aggregating local information from multi-level feature maps. The Dual Low-Rank Adaptation (Dual-LoRA) module projects the input feature into two low-rank subspaces: one for stable holistic domain knowledge (skill space) learning and the other for instruction condition (task space) learning. Dual-LoRA modules is integrated into the LLM’s query and value projection layers for efficiency.
@@ 3.1 Preliminaries
where $z_{i}$ is the output feature passed to the next layer, and $\alpha$ is a scaling factor. For convenience, we omit the indices $i$ in the following. Incorporating LoRA, the trainable parameters of $F$ are reduced to $\{\theta_{P},A,B\}$.
@@ 3.2 Framework Overview
A visual overview is provided in Fig.3. The proposed architecture introduces two key components: 1) a Visual Cue Enhancement (VCE) module, detailed in Sec.3.3, which enhances vision feature projection by aggregating multi-level local visual features, and 2) Dual Low-Rank Adaptation (Dual-LoRA), described in Sec. 3.4, which employs dual low-rank adaptation pathways to facilitate holistic-to-localized knowledge adaptation.
@@ 3.3 Multi-level Local Visual Cue Enhancement
Typical MLLMs rely on high-level vision feature maps to align vision and language modalities, e.g. LLaVA [20] and Qwen-VL [3] use the penultimate layer vision feature map of ViT [28]. While this map primarily captures high-level semantic features and may overlook local visual details. To address this limitation, we propose a lightweight Visual Cue Enhancement (VCE) module. This module enhances the projected feature map, i.e., the high-level feature $*F^{\text{*}}$, by integrating local information from other layers.
Let $[F_{l}]_{l=1}^{L}$ denote multi-level vision feature maps extracted from selected intermediate layers, where $l\in\{1,\ldots,L\}$ indexes the select layers. Suppose that the anchor patch feature to be enhanced is $F^{*}(p_{q})$, where $p_{q}\in\mathbb{R}^{2}$ is its 2D position. For each layer $l$, we employ deformable attention [34, 38] to extract local feature $F_{l}^{\prime}(p_{q})$:
$F_{l}^{\prime}(p_{q})$ $=\sum_{m=1}^{M}\sum_{k=1}^{K}A_{lmq}(k)\cdot W_{lm}F_{l}\left(p_{q}+\Delta p_{lmq}(k)\right),$ (5)
$A_{lmq}$ $=\mathrm{softmax}\left(\phi_{lm}^{a}(F_{l}(p_{q}))\right),$ (6)
$\Delta p_{lmq}$ $=\phi_{lm}^{o}(F_{l}(p_{q})),$ (7)
where $m\in\{1,\ldots,M\}$ indexes attention heads, $W_{lm}$ is a learnable projection weight for head $m$ at layer $l$, and $k\in\{1,\ldots,K\}$ indexes reference patches. The offset $\Delta p_{lmq}\in\mathbb{R}^{K\times 2}$ represents adaptively learned coordinate offsets around $p_{q}$, indicate the interest ares, while $A_{lmq}\in\mathbb{R}^{K}$ denotes attention weights for input $F_{l}(p_{q})$, where $\phi_{lm}^{o}$ and $\phi_{lm}^{a}$ are linear layers used for learning the offsets $\Delta p_{lmq}$ and attention $A_{lmq}$, respectively.
A Multi-Level Local Information Aggregation module is then used to consolidate local features across layers:
$F^{\prime}(p_{q})=W_{o}[\mathrm{concat}(F_{l}^{\prime}(p_{q}))_{l=1}^{L}],$ (8)
where $W_{o}$ projects the concatenated features into a unified space. Finally, the enhanced feature map is obtained through residual fusion and sent to the vision projector to generate enhanced vision tokens:
$t_{\mathrm{vision}}=P\left(\mathrm{Norm}\left(F^{*}+\gamma F^{\prime}\right);\theta_{P}\right),$ (9)
where $\gamma$ is a scaling parameter, and $\mathrm{Norm}(\cdot)$ denotes a LayerNorm [2] layer applied to normalize the residual.
@@ 3.4 Dual Low Rank Adaptation
In LoRA-MoE, the token embedding is passed through a router, typically a linear layer, followed by an activation strategy that determines which LoRA experts to activate. The activation strategies typically include sparse activation [5], dense activation [33], and rectified activation [13], each of which has distinct effects across various tasks.
However, efficiently balancing the activation strategy, the granularity of LoRA experts, and task complexity can be challenging. Moreover, the data passing through the router and multiple LoRA experts significantly increases inference time. For instance, a 4-expert LoRA-MoE uses 1.59 times the inference time compared to the non-MoE method (details can be found in Fig. 8). Hence, we aim to implement a unified approach that achieves the same effect as the LoRA-MoE method but more efficiently. We begin by analyzing the expressiveness of both LoRA-MoE and unified LoRA.
@@
The proof of Proposition 1 and its Corollary 1 is provided in Appendix A.
3.4.1 Learning Holistic Knowledge in One Skill Space
According to Corollary 1, a single LoRA with equivalent parameters has the same expressiveness as LoRA experts across various ranks. Let $\mathcal{S}$ denote the adaptation space, which we refer to as the skill space, modeled by the parameter matrix $S$.
In practice, however, LoRA-MoE often outperforms a single, larger LoRA [13]. What leads to the performance gap, despite they have the same expressiveness? This can be attributed to the differentiated local response capability of LoRA-MoE, i.e., expert activation, which provides redundant space to address data conflicts. Two methods ensure the local differences among experts are proposed: 1) variations in experts’ weight distributions at a given moment, e.g., RoDE [13] employs heterogeneous LoRA experts initialized with varying distributions, and 2) differences in routing scores among experts, which can lead to uneven expert activation, e.g., in extreme cases, only a single expert may be utilized [5].
The observation above suggests that $\mathcal{S}$ should respond locally (specifically) to the varying tasks. We now consider how to elegantly rectify $\mathcal{S}$ to achieve a “local response”.